# Why Isn’t Relational Learning Taking Over the World?

### v3

@@ Introduction
In my AI courses I often include the clicker question of Figure 1. Reading machine learning books and papers would lead one to assume the answer is A. Reading the press on generative AI would lead one to think the answer is B. Very few of the students choose A or B.
@@
    Entities and events (e.g., plants, people, diseases, lectures, university course)
@@
Learning and reasoning about entities, events and relations between them is central to AI:
@@
Generative AI is characterized by models that learn from huge datasets to predict probabilities from prompts and sampling from the distributions. For example, large language models can be seen as sampling from
$P(\mbox{\emph{Next, acc}}\mid\mbox{\emph{Con}})=P(\mbox{\emph{acc}}\mid\mbox{\emph{Next, Con}})P(\mbox{\emph{Next}}\mid\mbox{\emph{Con}})$
where Next is the next token, Con, the context, is the prompt and the previously generated text, and acc is the proposition that the text is acceptable from human feedback. $P(\mbox{\emph{Next}}\mid\mbox{\emph{Context}})$ is trained from enormous datasets. $P(\mbox{\emph{acc}}\mid\mbox{\emph{Next}},\mbox{\emph{Context}})$, reinforcement learning by human feedback (RLHF) [^1] To see how this is RLHF, $Acc$ is a reward of 1 for acceptable and 0 for unacceptable. Expected utility is equivalent to probability of Acc=1. Humans evaluate complete text. $P(\mbox{\emph{Acc}=1}\mid\mbox{\emph{Next}},\mbox{\emph{Context}})$ is the $Q$-function of reinforcement learning, where Next corresponds to the action and Context is the state. (Knox and Stone 2009; Ziegler et al. 2020), is learned from humans evaluating generated text.
@@
Contemporary relational learning tools are typically trained to predict probabilities, optimizing log loss. Evaluation is much more diverse; see below.
@@
The simplest form of relational data is exemplified by collaborative filtering datasets such as used in the Netflix challenge (Bell and Koren 2007) and Movielens datasets (Harper and Konstan 2015). Movielens contains rating datasets containing user-movie-rating-timestamp tuples, where the user and movie are identifiers represented by arbitrary numbers. There are also tables for user properties and movie properties, including links into the Internet Movie Database (IMDB). An even simpler version is a rated dataset that ignores the actual ratings, and is just a user-movie dataset (perhaps with timestamps). To predict rating (or rated) from just the rating dataset, you can define latent properties of users and of movies and use these to predict ratings. Representing latent properties as a vector, matrix factorization (Koren, Bell, and Volinsky 2009) learns about movies and users, without learning general knowledge. Another possible technique is to do a self-join, which can result in a user-user table or a movie-movie table.
@@
The second way to converting relations to triples is to use (row, column, value) triples. The row can be a primary key if there is one, or is a new identifier – the “reified” entity. Reify means “make into an entity” (or “thingify”). In Figure 2 (b), Q262802-3B1AE42E and Q262802-9c5c267f are reified entities. The number of facts about a reified entity is the number of columns in the table, which is typically just a handful. To convert to the original relation, one can do a database join of the triples on the reified entities and project onto the other properties (which is how I generated the table from the Wikidata triples). Reification does not lose information, but make the representation more modular: other properties, including meta-properties (such as the author) can be easily added to a knowledge graph.
@@
Big data often implies more entities, but less data, on average, about each entity. As entities are added, they start off as stubs with few tuples, resulting in a long tail of entities about which not much is known. As more detail is added about an entity, there are more associated reified entities, about which there are inherently few tuples (see Appendix A).
@@
Similar problems with standard benchmarks arise in the graph learning community (Bechler-Speicher et al. 2025; Roy et al. 2025). Although relational learning can be seen as an instance of graph learning, and graph learning an instance of relational learning, the challenges in this paper do not seem to have much in common with generic graph tasks.
@@
In the complete knowledge or closed world assumption, statements not in a database are false. E.g., Airlines know the actual departure and arrival times of flights in the past, who is booked on a flight, and who showed up for boarding. They do not know who will show up tomorrow. When data is considered complete up to some time, the predictive problem is to predict the future from the past. Consider, for example, the problem of predicting the passenger load or whether I will miss a connection, of a trip from Vancouver BC to Boston MA on July 2 2026. This is a tricky prediction as July 2 is a special day – is one day after a holiday in Canada, and 2 days before one in USA – and 2026 is different in that many Canadians are boycotting USA and the FIFA World Cup is being held then; Vancouver and Boston are host cities. Rather than guessing, we should be able to make an informed prediction based on data.
Surprisingly, there is very little work on making predictions about the future from arbitrary complete databases. We can’t add an airline reservation system to a learner and get it to predict future demand. It might be because most company databases are proprietary and include much metadata, such as the database schema. Individual companies such as Amazon and Netflix have huge datasets they are mining, but the techniques are proprietary and are undoubtedly tuned to the particular task, such as maximizing profit and engagement, rather than allowing general queries of what is predicted. Steck et al. (2021) describe techniques used by Netflix. Feature engineering (such as carefully chosen joins and projections) can created tables that methods such as XGBoost work on.
@@
Predicting property values, when the set of values is a constant set, such as the positions of football players, may be amenable to techniques for tabular learning, such as gradient tree boosting, following feature engineering. However, when making predictions where the values are entities (link prediction) this is infeasible.
@@
A standard to make probabilistic predictions is to use model with softmax or sigmoid output and to optimize log loss (an observation with predicted probability $p$ incurs a cost of $\log p$, the expectation of which is cross entropy) on the training set with some regularization.
@@
Ranking has many problems, including: following.
@@
Aggregation (e.g., predicting gender of a user from the movies they watched) is the Achilles heel of relational learning. Most models either have a built-in aggregation such as noisy-or for probabilistic logic programs (Poole 1993) or logistic regression for weighted logical formula models (Domingos and Lowd 2009; Kazemi et al. 2014), use a distribution of related entities in relational attention (Velickovic et al. 2018; Wang et al. 2019) or use explicit operations such a sum or mean (Schlichtkrull et al. 2018). They are problematic theoretically (e.g., asymptoticly) (Poole et al. 2014; Buchman and Poole 2015; Adam-Day et al. 2024) and in practice (Kazemi et al. 2017). They need to work with no related entities and as the number of related entities goes to infinity. The models either implicitly assume the related entities provide independent evidence (as do sum, noisy-or and logistic regression), or act the same as if there was just one related entity (as do max, average and attention). Determining whether the evidence is independent (so can be accumulated) or is dependent (in which case extra evidence can be discounted) is difficult, and invariably ignored.
@@ Appendix A Appendix A: Proportion of Wikidata entities with few tuples
The aim of this appendix is to estimate the proportion of entities in Wikidata that are the subject of few tuples. There is very long tail of entities about which there are few tuples, and presumably are stubs that will be expanded in the future, so it may be unfair to count those. Instead, we approximate the number of reified entities per top-level entity. I could not download all of Wikidata, so as a sample I selected the entities of the Wikipedia featured articles of March 2025 (https://en.wikipedia.org/wiki/Wikipedia:Today%27s˙featured˙article/March˙2025) that had Wikidata pages. The entities were diverse including Hughie Ferguson, a Scottish footballer in 1910s and 20s, Pulgasari, a North Korean film from 1985, the contemporary ballet Flight Pattern, Steele’s Greenville expedition in the American Civil War, the sun, and a fictional character in the medical drama series House.
For each entity (including reified entities), Wikidata allows users to download triples about that entity. For reified entities, the triples downloaded are the triples of the subject of the reification. We define the local entities for entity $e$ to be those that appear as a subject in a triple for the page for $e$ and for which Wikidata downloads the page for $e$. For each of the selected entities, we counted the number of local entities that are the subject of fewer than 10 triples and the number of entities that are the subject of 10 or more triples. See Figure 4. The number ranged from 5 for History of infant schools in Great Britain and 6 for Your Girl (a 2005 song by Mariah Carey), to 841 for the Sun and 520 for the composer and conductor Pierre Boulez. The entities with the lowest number of tuples are just stubs; note that for Q133116210, the 11 tuples included the name and labels in both English and Portuguese. The average number of local entities was 134 (or 108 if the top and bottom outliers were removed).
@@
Q3142353 125 50 5 Hughie Ferguson
Q931229 87 38 1 USS Congress
Q11813 1074 490 10 James Madison
Q3830799 161 63 1 Leroy Chollet
Q6529597 116 49 1 Les Holden
Q133116210 11 5 1 History of infant schools in Great Britain
Q4767002 210 89 1 Anna Filosofova
Q485172 684 285 6 Ann Arbor
Q16267674 426 282 3 Hotline Miami 2: Wrong Number
Q1349580 201 95 2 Matthew Brettingham
Q632511 325 113 4 Seattle Sounders FC
Q1137583 242 121 1 Pulgasari
@@
Q525 2605 841 2 Sun
Q122330832 66 47 2 All-American Bitch
Q130337417 32 17 1 The True Record
Q3856283 70 30 1 Michael Tritter
Q556464 270 158 1 Gertie the Dinosaur
Q5461994 201 112 1 Flotilla
Q156193 859 520 5 Pierre Boulez
Q545151 141 51 2 The Spy Who Loved Me
Q830590 62 32 1 Interstate 182
Q3142456 78 41 1 Hurricane Cindy
@@
Figure 4: For each top-level entity: the Wikidata ID, the number of tuples with the entity as subject, the number of local entities that are the subject of fewer than 10 tuples, the number with 10 or more tuples, and the English name of the entity. These numbers were from 2025-08-19; they change over time as the knowledge base evolves. The Python code to generate the contents of this table is available from the author.
@@ Appendix B Appendix B: An MRR agent
Companies that present lists of results and get feedback about which item(s) was selected or how many items were presented, such as Google, will have data on the distribution of the number of items considered. It would be better to use this distribution for evaluation than MRR, but it probably depends on what is being presented (and whether predictions can be tested to determine if that are correct or what is wanted).

### v4

@@ Introduction
In my AI courses, I often include the clicker question of Figure 1. Reading machine learning books and papers would lead one to assume the answer is A. Reading the press on generative AI would lead one to think the answer is B. Very few of the students choose A or B.
@@
    Entities and events (e.g., plants, people, diseases, lectures, university courses)
@@
Learning and reasoning about entities, events and relations among them is central to AI:
@@
The simplest form of relational data is exemplified by collaborative filtering datasets such as used in the Netflix challenge (Bell and Koren 2007) and Movielens datasets (Harper and Konstan 2015). Movielens contains rating datasets containing user-movie-rating-timestamp tuples, where the user and movie are identifiers represented by arbitrary numbers. There are also tables for user properties and movie properties, including links into the Internet Movie Database (IMDB). An even simpler version is a rated dataset that ignores the actual ratings, and is just a user-movie dataset (perhaps with timestamps). To predict rating (or rated) from just the rating dataset, you can use latent properties (embeddings) of users and of movies to predict ratings. Representing latent properties as a vector, matrix factorization (Koren, Bell, and Volinsky 2009) learns about movies and users, without learning general knowledge. Another possible technique is to do a self-join, which can result in a user-user table or a movie-movie table.
@@
The second way to convert relations to triples is to use (row, column, value) triples. The row can be a primary key if there is one, or is a new identifier – the “reified” entity. Reify means “make into an entity” (or “thingify”). In Figure 2 (b), Q262802-3B1AE42E and Q262802-9c5c267f are reified entities. The number of facts about a reified entity is the number of columns in the table, which is typically just a handful. To convert to the original relation, one can do a database join of the triples on the reified entities and project onto the other properties (which is how I generated the table from the Wikidata triples). Reification does not lose information, but make the representation more modular: other properties, including meta-properties (such as the author) can be easily added to a knowledge graph.
@@
Note that graph learning (Hamilton 2020) and relational learning are isomorphic problems. A graph can be represented as a tail-head-label “arc” relation. Relations can be represented as knowledge graphs. These are complementary fields. It is not clear that a typical relational database is anything like a typical graph. Knowledge graphs tend to be much more heterogenous than graph benchmarks (Bechler-Speicher et al. 2025); the only general structure is that induced by reification. Many of the questions of interest are different, for example, there tends to be only a single knowledge graph so graph matching (Roy et al. 2025) is not really a problem of interest. Most of the issues in this paper do not apply to graph learning in general.
@@
Big data often implies more entities, but less data, on average, about each entity. As entities are added, they start off as stubs with few tuples, resulting in a long tail of entities about which little is known. As more detail is added about an entity, there are more associated reified entities with inherently few tuples (see Appendix A).
@@
Graph learning also has issues with standard benchmarks (Bechler-Speicher et al. 2025; Roy et al. 2025).
@@
In the complete knowledge or closed world assumption, statements not in a database are false. E.g., Airlines know the actual departure and arrival times of flights in the past, who is booked on a flight, and who showed up for boarding. They do not know who will show up tomorrow. When data is considered complete up to some time, the predictive problem is to predict the future from the past. Consider, for example, the problem of predicting the passenger load or whether I will miss a connection, of a trip from Vancouver BC to Boston MA on July 2 2026. This is a tricky prediction as July 2 is a special day – is one day after a holiday in Canada, and 2 days before one in USA – and 2026 is unusual in that many Canadians are boycotting USA and the FIFA World Cup is being held then; Vancouver and Boston are host cities. Rather than guessing, we should be able to make an informed prediction based on data.
Surprisingly, there is very little work on making predictions about the future from arbitrary complete databases. We can’t add an airline reservation system to a learner and get it to predict future demand. It might be because most company databases are proprietary and include much metadata, such as the database schema. Individual companies such as Amazon and Netflix have huge datasets they are mining, but the techniques are proprietary and are undoubtedly tuned to the particular task, such as maximizing profit and engagement, rather than allowing general queries of what is predicted. Steck et al. (2021) describe techniques used by Netflix. Feature engineering (such as carefully chosen joins and projections) can created tables that methods such as XGBoost or deep learning can work on.
@@
Predicting property values, when the set of values is a constant set, such as the positions of football players, may be amenable to techniques for tabular learning, such as gradient tree boosting, following feature engineering. However, when making predictions where the values are entities (called link prediction) this is infeasible.
@@
A standard way to make probabilistic predictions is to use a softmax or sigmoid output and to optimize log loss (an observation with predicted probability $p$ incurs a cost of $\log p$, the expectation of which is cross entropy) on the training set with some regularization.
@@
When considering relational data as a graph, the neighbours of an entity are typically either all entities or those known to be related by existing relations. When considering entities $n$ hops away from an entity, the number of entities is exponential in $n$. This is, of course, bounded by the number of entities. Having all entities as neighbours of an entity allows for richer hypotheses about which pairs are related, but impacts on what methods are appropriate. With large knowledge bases, algorithms that are quadratic in the number of entities are impractical. Learning needs to be done in linear time, at most, and inference in sublinear time. Fortunately many questions can be approximated by sampling.
@@
Ranking has many problems, including:
@@
Aggregation (e.g., predicting gender of a user from the movies they watched) is the Achilles heel of relational learning. Most models either have a built-in aggregation such as noisy-or for probabilistic logic programs (Poole 1993) or logistic regression for weighted logical formula models (Domingos and Lowd 2009; Kazemi et al. 2014), use a distribution of related entities in relational attention (Velickovic et al. 2018; Wang et al. 2019) or use explicit operations such a sum or mean (Schlichtkrull et al. 2018). They are problematic theoretically (e.g., asymptoticly) (Poole et al. 2014; Buchman and Poole 2015; Adam-Day et al. 2024) and in practice (Kazemi et al. 2017). They need to work with no related entities and as the number of related entities goes to infinity. The models either implicitly assume the related entities provide independent evidence (as do sum, noisy-or and logistic regression), or act the same as if there were one or few related entities (as do max, average and attention). Determining whether the evidence is independent (so can be accumulated) or is dependent (in which case extra evidence can be discounted) is difficult, and invariably ignored.
@@
Acknowledgements
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC). Thanks to Sriraam Natarajan for comments on an earlier version.
@@ Appendix A Appendix A: Proportion of Wikidata entities with few tuples
The aim of this appendix is to estimate the proportion of entities in Wikidata that are the subject of few tuples. There is very long tail of entities about which there are few tuples, and presumably are stubs that will be expanded in the future, so it may be unfair to count those. Instead, I approximate the number of reified entities per top-level entity. I could not download all of Wikidata, so as a sample I selected the entities of the Wikipedia featured articles of March 2025 (https://en.wikipedia.org/wiki/Wikipedia:Today%27s˙featured˙article/March˙2025) that had Wikidata pages. The entities were diverse including Hughie Ferguson, a Scottish footballer in 1910s and 20s, Pulgasari, a North Korean film from 1985, the contemporary ballet Flight Pattern, Steele’s Greenville expedition in the American Civil War, the sun, and a fictional character in the medical drama series House.
For each entity (including reified entities), Wikidata allows users to download triples about that entity. For reified entities, the triples downloaded are the triples of the subject of the reification. Define the local entities for entity $e$ to be those that appear as a subject in a triple for the page for $e$ and for which Wikidata downloads the page for $e$. For each of the selected entities, I counted the number of local entities that are the subject of fewer than 10 triples and the number of entities that are the subject of 10 or more triples. See Figure 4. The number ranged from 5 for History of infant schools in Great Britain and 6 for Your Girl (a 2005 song by Mariah Carey), to 841 for the Sun and 520 for the composer and conductor Pierre Boulez. The entities with the lowest number of tuples are just stubs; note that for Q133116210, the 11 tuples included the name and labels in both English and Portuguese. The average number of local entities was 132 (or 108 if the top and bottom outliers were removed).
@@
Q525 2605 841 2 Sun
Q11813 1074 490 10 James Madison
Q156193 859 520 5 Pierre Boulez
Q485172 684 285 6 Ann Arbor
Q192920 475 172 2 Edward the Martyr
Q16267674 426 282 3 Hotline Miami 2: Wrong Number
Q632511 325 113 4 Seattle Sounders FC
Q556464 270 158 1 Gertie the Dinosaur
Q1137583 242 121 1 Pulgasari
Q822900 214 36 1 geography of Ireland
Q4767002 210 89 1 Anna Filosofova
Q1349580 201 95 2 Matthew Brettingham
Q5461994 201 112 1 Flotilla
Q3830799 161 63 1 Leroy Chollet
Q545151 141 51 2 The Spy Who Loved Me
Q3142353 125 50 5 Hughie Ferguson
Q6529597 116 49 1 Les Holden
Q931229 87 38 1 USS Congress
Q3142456 78 41 1 Hurricane Cindy
@@
Q3856283 70 30 1 Michael Tritter
Q122330832 66 47 2 All-American Bitch
Q830590 62 32 1 Interstate 182
Q130337417 32 17 1 The True Record
@@
Q133116210 11 5 1 History of infant schools in Great Britain
@@
Figure 4: For each top-level entity: the Wikidata ID, the number of tuples with the entity as subject, the number of local entities that are the subject of fewer than 10 tuples, the number with 10 or more tuples, and the English name of the entity. Sorted by number of tuples. These numbers were from 2025-08-19; they change over time as the knowledge base evolves. The Python code to generate the contents of this table is available from the author.
@@ Appendix B Appendix B: An MRR agent
Companies that present lists of results and get feedback about which item(s) was selected or how many items were presented, such as Google, will have data on the distribution of the number of items considered. It would be better to use this distribution for evaluation than MRR, but it probably depends on what is being presented (and whether predictions can be tested to determine if they are correct or what is wanted).